### v1

# CleanSurvival: Automated data preprocessing for time-to-event models using reinforcement learning

Abstract
Data preprocessing is a critical yet frequently neglected aspect of machine learning, often paid little attention despite its potentially significant impact on model performance. While automated machine learning pipelines are starting to recognize and integrate data preprocessing into their solutions for classification and regression tasks, this integration is lacking for more specialized tasks like survival or time-to-event models. As a result, survival analysis not only faces the general challenges of data preprocessing but also suffers from the lack of tailored, automated solutions in this area.
To address this gap, this paper presents CleanSurvival, a reinforcement-learning-based solution for optimizing preprocessing pipelines, extended specifically for survival analysis. The framework can handle continuous and categorical variables, using $Q$-learning to select which combination of data imputation, outlier detection and feature extraction techniques achieves optimal performance for a Cox, random forest, neural network or user-supplied time-to-event model. The package is available on GitHub: https://github.com/datasciapps/CleanSurvival.
Experimental benchmarks on real-world datasets show that the $Q$-learning-based data preprocessing results in superior predictive performance to standard approaches, finding such a model up to 10 times faster than undirected random grid search. Furthermore, a simulation study demonstrates the effectiveness in different types and levels of missingness and noise in the data.
1 Introduction
In the era of big data and machine learning (ML), the ability to extract meaningful insights from complex datasets is paramount. A critical step in this process is data preprocessing, which involves cleaning, transforming, and preparing raw data to be suitable for analysis. The quality of data preprocessing may significantly impact the performance and reliability of ML models. This is particularly crucial in the field of survival analysis, where the goal is to predict the time until an event of interest occurs, such as patient death, equipment failure or customer churn.
Survival analysis poses unique challenges due to the presence of censored data, where the event of interest has not yet been observed. However, it is also overlooked in the context of automated machine learning (AutoML) pipelines, which aim to streamline the ML development process by automating tasks such as algorithm selection, hyperparameter tuning and model evaluation—and increasingly incorporate data preparation as part of their pipelines. In this paper, we introduce CleanSurvival, an automated data preprocessing framework tailored for survival analysis.
CleanSurvival leverages reinforcement learning (RL) techniques, specifically $Q$-learning, to optimize decisions such as imputation of missing values, detection and handling of outliers and feature extraction for survival models. RL is a powerful approach for automated data preprocessing because it dynamically optimizes pipeline steps by learning to maximize a reward function tied directly to model performance, ensuring data cleaning decisions are guided by their impact on survival predictions. The framework is designed to handle continuous and categorical variables, and can be used with a variety of time-to-event models, from classical methods to modern deep learning frameworks.
The framework, available as an open-source Python package, is demonstrated on several common survival analysis datasets, highlighting the sensitivity of predictive performance to data preprocessing steps and boasting improved predictive performance compared to standard approaches.
The article is organized as follows. Section 2 provides an overview of data preprocessing, survival analysis, AutoML and $Q$-learning. Section 3 reviews existing approaches to automated data preprocessing and AutoML frameworks. Section 4 describes the architecture of CleanSurvival and its features. Section 5 presents the results of experimental evaluation of the framework on real-world datasets. Finally, Section 6 discusses the results and outlines future directions for research.
2 Background
Data preprocessing involves cleaning, transforming and organizing raw data into a suitable format for analysis, and is a important step in the ML pipeline. Sub-tasks of data preprocessing include imputation or removal of missing values, detection and handling of outliers, variable selection and feature extraction and data transformations. These steps can have a profound downstream impact on classification performance [15] and model explanations [30].
However, the selection of appropriate preprocessing methods often requires a combination of domain knowledge, visual inspection and manual experimentation; it is also often poorly documented, whether in academic papers or computational notebooks [32, 9]. Some authors have even attempted to quantify the effect of preprocessing steps on model predictions independently of the dataset [10].
Automated data preprocessing has emerged to address these challenges [3, 28, 27, 19]. This approach uses algorithms and heuristics to automate various data cleaning and transformation tasks, reducing the need for manual intervention or iteration and potentially improving the efficiency and effectiveness of the preprocessing stage, ideally by learning from past cleaning tasks [18]. However, the field is still in its infancy.
2.1 Survival analysis
Survival analysis, also known as reliability analysis or duration modelling, is a statistical method for analysing time-to event data. It is widely used in various fields, including medicine, engineering and social sciences. In survival analysis, the primary goal is to model the time until an event of interest occurs, such as death, disease progression, machine failure or customer churn. The survival function,
$S(t)=P(T>t),$
denotes the probability that the time of event (death), a random variable $T$, occurred later than a time $t$. A unique characteristic of time-to-event problems is censoring, or data points that are only partially observed, such as patients who survived up until the last observation time, at which point they were lost to follow-up or the study period ended [36].
Naïvely, one can treat survival analysis as either a regression or classification problem, but both approaches lead to a significant information loss. In the former case, one treats the observed survival time as a continuous outcome, and censored observations are either discarded or imputed, resulting in substantial reduction in sample size or bias introduced by oversimplified assumptions about the censoring process. In the latter case, one models survival—or not—in a predefined time window, reducing the problem to binary classification and losing granular information about event times [14].
Survival analysis models, such as the Cox proportional hazards model, Kaplan–Meier estimator and accelerated failure time models, are widely used in practice. These models estimate the hazard function, survival function, or survival probabilities over time, providing valuable insights into the relationship between covariates and survival outcomes.
Concordance indices like the C-index are widely used in evaluation of survival models due to their simplicity and ease of interpretation [16]. The C-index evaluates the model’s ability to correctly rank individuals based on their risk of experiencing the event. A higher C-index indicates better discriminatory power.
However, concordance indices have significant limitations, due to their poor calibration and failure to consider the distribution of survival times as well as their ranks. To measure how well survival probabilities align with observed event times, calibration metrics like Houwelingen’s $\alpha$ or D-calibration can be used [35] , and should be compared against a baseline model, such as the Kaplan–Meier estimator. Weighted integrated survival log loss or integrated Graf score are recommended scoring rules [31].
To account for censoring, we apply inverse probability of censoring weighting (IPCW). Let $G(t)$ be the Kaplan–Meier estimate of the probability of not being censored at $t$. Then, the integrated Graf score is defined
$\text{IGS}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{k}\Delta t_{j}\,\frac{\left(S(t_{j}|x_{i})-\text{Observed}(t_{j},i)\right)^{2}}{G(t_{j})},$
for a set of time points $\{t_{1},t_{2},\ldots,t_{k}\}$, where $\Delta t_{j}$ represents the difference between time points, $n$ is the number of individuals and $\text{Observed}(t,i)$ is an indicator function, equal to $1$ if the individual is known to have survived beyond $t$ and 00 otherwise.
2.2 Automated machine learning
Automated machine learning (AutoML) aims to streamline the process of ML development by automating steps such as algorithm selection, hyperparameter tuning and model evaluation, reducing the amount of time and expertise required by practitioners to train, deploy and fine-tune models [8, 1]. AutoML frameworks have seen success in various applications by employing search strategies such as meta-learning, Bayesian optimization and ensemble learning to achieve competitive performance.
However, data preprocessing remains an important analysis step that typically falls outside the AutoML pipeline [23]. Data preparation steps including cleaning, normalization and feature engineering are critical for the success of ML models but can be highly problem-specific [12]. Automating these tasks while maintaining flexibility for diverse datasets remains a significant hurdle [27, 19]. Mahdavi et al. [18] highlighted the potential of AI to solve data quality problems through data profiling and learning from past cleaning attempts [see also 17]. A holistic approach integrates the cleaning process with downstream tasks so that the cleaning is optimized for predictive performance [20]; indeed when integrated into an AutoML framework some cleaning steps may be more important than others [21].
Survival analysis in particular faces particular challenges, partly due to the relative lack of support for such models in the first place (versus classification or regression), as well as unique difficulties of handling censored time-to-event data [36], which are typically not addressed in conventional AutoML frameworks and do not feature in AutoML surveys [e.g. 1]. Seamlessly integrating these domain-specific preprocessing steps with the downstream tasks of model optimization and evaluation is a complex, underexplored area.
Prominent AutoML pipelines include Auto-WEKA, an early AutoML system that uses Bayesian optimization to search for the best combination of preprocessing steps and machine learning algorithms [34]; TPOT, a tree-based pipeline optimization tool that uses genetic programming to evolve pipelines of data cleaning and machine learning operations [22]; and auto-sklearn, an extension of Auto-WEKA that incorporates more recent advancements in machine learning and hyperparameter optimization while offering a familiar interface based on the Python package scikit-learn [8].
These AutoML pipelines have demonstrated promising results in various domains, including image classification, natural language processing and tabular data analysis. However, they often focus on general machine learning tasks and may not be specifically tailored to the challenges of survival analysis, particularly in the presence of missing data.
Salhi et al. [27] presented a recent survey of data preprocessing using AutoML (though survival analysis is not mentioned). In their review, they highlight the relative capabilities of AutoML platforms: in many cases the data processing support is relatively basic. The authors indicate that all 11 tools reviewed support missing value imputation, but also state: ‘auto-sklearn cannot handle missing values’ and this must be done manually by the user—a contradiction. In this case, claims of support for data processing may actually be based on those of the underlying, non-automated ML framework (i.e. scikit-learn). Mumuni and Mumuni [19] also surveyed automated data processing for deep learning applications, highlighting in more detail the extent to which data processing steps are integral components of the automated pipeline. They similarly note the lack of early support from autosklearn.
2.3 $Q$-learning
Reinforcement learning is well-suited to the task of automating constrained ML pipelines, as it optimizes sequential decision-making processes, balancing exploration and exploitation, while being less computationally intensive than other methods, such as unconstrained evolutionary algorithms [11].
$Q$-learning is a model-free off-policy reinforcement learning algorithm that seeks to find the optimal action-selection policy for an agent interacting with an environment. The algorithm is based on estimating the value $Q=Q(s,a)$ of taking an action $a$ in a given state $s$. The agent iteratively updates these $Q$ values based on its experiences, enabling it to learn an optimal policy even in environments with stochastic rewards and transitions.
The goal of $Q$-learning is to maximize the cumulative reward over time by updating $Q$ according to the Bellman equation. Given a current state $s$, action $a$, reward $r$ and next state $s^{\prime}$, the update rule is:
${Q(s,a)\leftarrow Q(s,a)+\alpha\bigl{(}r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\bigr{)},}$ (1)
where $\alpha$ is the learning rate and $\gamma$ is the discount factor, controlling the importance of future rewards. The update equation Equation 1 allows the $Q$-learning agent to converge to an optimal policy $\pi^{*}$, defined
${\pi^{*}(s)=\arg\max_{a}Q(s,a).}$ (2)
without requiring a model of the environment’s dynamics. By exploring various state–action pairs and refining $Q$-values, $Q$-learning is able to asymptotically approach optimal behaviour, provided that the agent balances exploration and exploitation effectively.
Alternatives to $Q$-learning, such as Bayesian optimization, can offer improved sample efficiency in some cases but often struggle to scale in high-dimensional or discrete search spaces typical of complex AutoML pipelines. More specialized methods, such as neural architecture search, are not easily extensible to data preprocessing tasks. Other reinforcement learning approaches, including deep reinforcement learning [11] and Monte Carlo tree search [7], provide flexibility but introduce additional computational overhead and may require constraints or tailored mechanisms to ensure valid ML pipelines.
3 Related work
Table 1 compares the features of various AutoML solutions; as highlighted in Section 2, few offer integrated support for survival analysis and data preprocessing is not always within the optimization loop.
Table 1: Comparison of AutoML frameworks and their native support for survival analysis and dynamic data preprocessing
Framework Method Preprocessing Survival
Amazon SageMaker Autopilot Bayesian optimization and ensembles Yes No
auto-keras Neural architecture search Yes No
auto-sklearn Bayesian optimization Limited No
AutoGluon Stack ensembling Yes No
Azure AutoML Bayesian optimization and meta-learning Yes No
BigML Decision tree-based optimization Yes No
DataRobot Proprietary ensemble and optimization Yes Limited
FLAML Cost-aware Bayesian optimization No No
Google AutoML Tables Neural architecture search Yes No
H2O AutoML Random search and stacked ensembles Yes No
MLflow Manual configuration Yes No
MLJAR Random search and stacked ensembles Yes No
PyCaret Iterative search with pipeline tuning Yes Limited
TPOT Genetic programming No No
Of frameworks offering automated data cleaning, DataRobot is proprietary, commercial platform and only appears to offer time-to-event modelling via a ‘hack’ of converting the task to a classification problem via discretization [see e.g. 5]. Meanwhile, H2O.ai can run Cox proportional hazards models as a fixed model, but not via its AutoML interface.
However, some dedicated data cleaning solutions have been proposed. Bilal et al. [3] proposed Auto-Prep, an interactive Python-based tool that recommends data cleaning methods to the user based on application of candidate techniques and subsequent evaluation using simple classifiers or regression models. In a review of data preprocessing in AutoML (Section C) the authors highlight the capabilities, or lack thereof, of AutoML tools to perform data preprocessing and feature engineering without manual human intervention. Another Python package, Atlantic [28], automates preprocessing steps including feature engineering and missing value imputation for supervised learning tasks. The framework identifies the best combination of steps based on evaluation using tree-based model ensembles.
Berti-Equille [2] developed Learn2Clean, a tool offering an innovative approach to data preprocessing. It leverages $Q$-Learning, a reinforcement learning technique, to dynamically select the optimal sequence of preprocessing tasks for a given dataset and ML model. This optimization aims to maximize the quality of the ML model’s results. Learn2Clean implements automated data preprocessing for regression, classification and clustering tasks, using $Q$-learning to optimize respective evaluation criteria: mean squared error, accuracy and silhouette index. However, Learn2Clean limitations include lack of built-in support for survival analysis, categorical data types, flexible hyperparameter tuning and custom reward functions. It also has a complex dependency structure, which can make initial setup challenging for end-users.
MLsurvival [37] described an automated tool for cancer survival prediction that removes or imputes missing values, selects and standardizes features, trains survival models and then makes predictions. Unfortunately, neither a full text article nor open source implementation of the method were published. More recently, Pomsuwan and Freitas [25] proposed an AutoML system for survival analysis based on genetic algorithms and a combination of elastic-net Cox models, random survival forests and survival trees, optimizing for C-index. However, the tool does not incorporate data preprocessing.
4 Methodology
In this section we describe CleanSurvival, our proposed AutoML data preprocessing tool for survival analysis, illustrated in Figure 1.
Figure 1: Architecture of the CleanSurvival automated data cleaning framework
4.1 Data preprocessing methods
4.1.1 Missing values
CleanSurvival offers a variety of methods for handling missing values, addressing different data characteristics and analytical goals.
  • For straightforward scenarios with minimal missingness, complete case analysis (CCA) simply removes rows containing missing values.
  • Simple mean/median imputation replaces missing values with the mean or median of the observed values for a given variable.
  • Multiple imputation using chained equation [MICE, 4] provides a more robust approach by generating multiple suitable replacements for each missing value, creating several complete datasets for analysis. This method utilizes an iterative imputer, which starts with an initial guess and refines the estimates until convergence.
  • Finally (but most computationally intensive): $k$-nearest neighbors (KNN) imputation identifies the $k$ most similar observations to the one with missing values, based on other, non-missing features, and uses their values to impute the missing data. Mean or median imputation is a limiting case where $k\rightarrow n$.
4.1.2 Outliers
To ensure the reliability and validity of survival analysis, robust outlier detection methods are essential.
  • For multivariate datasets, CleanSurvival employs the Elliptic Envelope algorithm [26] to identify and remove outliers based on their Mahalanobis distances from the data centre. This method is particularly useful for detecting outliers that deviate from the overall correlation structure of the data.
  • The Martingale residuals method [33] calculates the difference between the observed and expected number of events for each individual (based on a simple Kaplan–Meier estimator), providing a measure of how unusual their survival time is compared to the expected survival time.
4.1.3 Variable selection and feature extraction
To identify the most salient variables for survival analysis, a variety of feature selection methods are available, enhancing both model performance and interpretability.
  • The Univariate Cox Proportional Hazards Selection (UC) method assesses the individual effect of each feature on survival using the Cox Proportional Hazards model. It selects features based on the significance of their coefficients, highlighting variables strongly associated with survival outcomes.
  • The LASSO (Least Absolute Shrinkage and Selection Operator) regression technique shrinks the coefficients of less important features to zero, effectively performing feature selection.
  • Recursive Feature Elimination (RFE) recursively removes the least important features based on their contribution to a model’s performance, using cross-validation to evaluate the model’s performance at each step.
  • The Information Gain Selection (IG) method measures the amount of information gained about the target variable (survival outcome) by knowing the value of a feature. This helps identify the most relevant variables by selecting features that provide the most information about the survival outcome.
4.2 Survival analysis
Three survival analysis models were carefully selected to integrate into CleanSurvival, each chosen for its unique strengths and applicability to a variety of survival analysis scenarios.
Cox proportional hazards
This widely-used model [6] is valued for its interpretability, allowing researchers to quantify the impact of different factors on the hazard rate. Its assumption of proportional hazards can be a limitation in some cases.
Random survival forest
The RSF model [24] is an ensemble method based on decision trees, offering robustness to nonlinearities and interactions in the data. Its non-parametric nature makes it a flexible choice when the underlying relationships in the data are not well understood.
DeepHit neural network
This deep learning model [13] leverages the power of neural networks to capture complex patterns and interactions in survival data. Its ability to model multiple competing risks makes it particularly well-suited for scenarios where individuals may experience different types of events.
4.2.1 Reward structure
To guide the $Q$-learner effectively for survival analysis problems, the reward structure is adapted to use the concordance index or C-index [16].
4.3 Working modes
To provide users with a range of options for data preprocessing and analysis, four distinct working modes were implemented in CleanSurvival:
Main algorithm
This mode uses the core $Q$-learning algorithm to identify the optimal sequence of preprocessing steps that maximize the performance of the chosen survival analysis model. This is the primary mode for automated pipeline optimization.
Random cleaning
In this mode, users can specify the desired number of random experiments. The tool generates random preprocessing pipelines and evaluates their performance, providing insights into the impact of different preprocessing choices. This mdoe can serve as a baseline for comparison with the optimized pipeline.
Custom pipeline
This mode allows users to define their own fixed preprocessing pipelines using a simple text configuration file. Each line in the file specifies a sequence of preprocessing methods, providing flexibility for testing specific hypotheses or domain knowledge.
No preparation
This mode bypasses all preprocessing steps, directly passing the raw dataset to the chosen survival analysis model. This can be useful for establishing a baseline for performance without any preprocessing.
The inclusion of these working modes significantly enhances the utility of the framework by offering valuable baselines for evaluating the effectiveness of the optimized pipelines generated by the $Q$-learning algorithm.
5 Experiments
5.1 Experimental setup
All methods were implemented in Python. Source code and documentation are available at https://github.com/datasciapps/CleanSurvival. Details of the datasets used are provided in Table 2. We demonstrate the approach using the rotterdam dataset, derived from the Rotterdam Study, a large prospective cohort study in the Netherlands, and flchain, a study of the relationship between serum free light chain (a type of blood measurement) and mortality.
Table 2: Summary of datasets used in the experiments
Dataset Samples Features Source
Rotterdam 2982 14 Rotterdam Study
Flchain 7874 11 UCI Machine Learning Repository
The results of data preprocessing strategies suggested by CleanSurvival are compared against the following methods:
  1. 1.
    complete case analysis (CCA), effectively ignoring the problem of missingness and outliers
  2. 2.
    random selection of imputation methods, simulating a grid search over possible analysis choices
  3. 3.
    mean imputation, as an example of a reasonable baseline used by an analyst not exploring the sensitivity of the model to different imputation strategies
We evaluate each method based on quality metrics as well as the time taken to retrieve an acceptable solution.
Hyperparameters that are not part of the CleanSurvival search space were left at default values, with architectural choices for the deep learning framework as follows: the DeepHit architecture consists of a shared sub-network with 2 hidden layers, each containing 100 neurons, using ReLU as the activation function. This is followed by cause-specific sub-networks, each containing 2 hidden layers with 50 neurons per layer. Dropout regularization was applied to all layers with a keep probability of 0.8 to mitigate overfitting. The model was trained using the Adam optimizer with a learning rate of 0.001 for 10,000 iterations. In practice, these options are customizable by the user or could be incorporated into the reinforcement learning action space as additional modules.
Since the Rotterdam dataset does not contain missing values, three variations of the dataset were used each representing a different missingness strategy (MCAR, MAR, MNAR) which were generated using the Jenga framework [29].
5.2 Results
Table 3: Comparing CleanSurvival to complete case analysis and mean imputation
Dataset Missingness CleanSurvival CCA Mean
Rotterdam 50% MCAR 0.833 0.803 0.800
Rotterdam 50% MAR 0.835 0.815 0.797
Rotterdam 50% MNAR 0.833 0.778 0.795
Flchain Existing 0.695 0.621 0.655
(a) MCAR
(b) MAR
(c) MNAR
(d) Flchain
Figure 2: CleanSurvival vs Grid Search
The results illustrated in Figure 2 clearly demonstrate the efficiency of CleanSurvival in identifying optimal data preprocessing pipelines for survival analysis within a limited timeframe of 5 mins. It can be seen that, CleanSurvival reaches the optimal solution in 9-14% of the time taken by a bruteforce to reach an optimal solution, translating to a time reduction of approximately 86-91%. Compared to a traditional grid search approach, which explores all possible pipeline combinations, CleanSurvival consistently achieves comparable or superior performance in a fraction of the time—reducing computational time by up to 80% in most cases. This highlights the effectiveness of the reinforcement learning approach in navigating the complex search space of data preprocessing options and efficiently identifying the most impactful transformations.
Specifically, the graphs show that CleanSurvival reaches the optimal C-index value significantly faster than the grid search, often within the first 20-30 seconds of the experiment compared to several minutes for grid search approaches. This rapid convergence to the optimal solution is crucial in practical settings where time constraints are common. Moreover, even when both methods eventually reach similar C-index values, CleanSurvival does so with considerably less computational effort, as evidenced by the shorter time-to-optimal, which on average was reduced by approximately 83%. This efficiency gain can be attributed to the intelligent exploration and exploitation strategy employed by the reinforcement learning agent, which allows it to focus on the most promising areas of the search space and avoid unnecessary evaluations.
These findings highlight the potential of CleanSurvival as a valuable tool for accelerating and automating the data preprocessing stage of survival analysis. By efficiently identifying optimal pipelines, not only does it save time and resources but it also enables researchers to focus on the core aspects of their analysis, ultimately leading to more robust and reliable results.
6 Conclusion
In this paper, we have introduced CleanSurvival, the first ever $Q$-learning-based framework that automates data preprocessing for survival analysis, addressing the often underserved challenge of automation in presence of censored data. We demonstrated its ability to enhance predictive performance and computational efficiency across different settings, compared to conventional methods or heuristic approaches.
The framework’s adaptability to various survival analysis models and its support for diverse preprocessing techniques make it a valuable tool for researchers and practitioners. Experimental results confirm that CleanSurvival not only accelerates the discovery of optimal pipelines but also maintains robust performance under different missingness patterns. These findings underscore the critical role of automated data preprocessing in enhancing the reliability of survival models and the feasibility of integrating reinforcement learning techniques into AutoML workflows.
Future work will focus on extending the framework to handle additional preprocessing tasks, incorporating advanced reinforcement learning strategies and improving scalability for large datasets and complex pipelines. Additionally, ensembling over differently cleaned datasets, together with raw data, may have the potential to increase predictive performance.
7 Acknowledgement
We are grateful for the input of Sergey Redyuk, whose suggestions greatly improved the manuscript. This work was supported by the German Federal Ministry of Education and Research via the project “Eventful” under grant ID 01IW23005.

### v2

@@
 Abstract
-Data preprocessing is a critical yet frequently neglected aspect of machine learning, often paid little attention despite its potentially significant impact on model performance.
+Background: Data preprocessing is a critical yet frequently neglected aspect of machine learning, often paid little attention despite its potentially significant impact on model performance.
 While automated machine learning pipelines are starting to recognize and integrate data preprocessing into their solutions for classification and regression tasks, this integration is lacking for more specialized tasks like survival or time-to-event models.
 As a result, survival analysis not only faces the general challenges of data preprocessing but also suffers from the lack of tailored, automated solutions in this area.
-To address this gap, this paper presents CleanSurvival, a reinforcement-learning-based solution for optimizing preprocessing pipelines, extended specifically for survival analysis.
-The framework can handle continuous and categorical variables, using $Q$-learning to select which combination of data imputation, outlier detection and feature extraction techniques achieves optimal performance for a Cox, random forest, neural network or user-supplied time-to-event model.
-The package is available on GitHub: https://github.com/datasciapps/CleanSurvival.
-Experimental benchmarks on real-world datasets show that the $Q$-learning-based data preprocessing results in superior predictive performance to standard approaches, finding such a model up to 10 times faster than undirected random grid search.
+Method: To address this gap, this paper presents CleanSurvival, a reinforcement-learning-based solution for optimizing preprocessing pipelines, extended specifically for survival analysis.
+The framework can handle continuous and categorical variables.
+It builds upon Learn2Clean’s $Q$-learning to select which combination of data imputation, outlier detection and feature extraction techniques achieves optimal performance for a Cox, random forest, neural network or user-supplied time-to-event model.
+The package is available on GitHub: https://github.com/phoenix0401/CleanSurvival.
+Results: Experimental benchmarks on real-world datasets show that the $Q$-learning-based data preprocessing results in superior predictive performance to standard approaches, finding such a model up to 10 times faster than undirected random grid search.
 Furthermore, a simulation study demonstrates the effectiveness in different types and levels of missingness and noise in the data.
+Conclusion: With an increase in the use of machine learning, it becomes important to generalise AutoML pipelines to a variety of models now present, including survival analysis.
+Tools like CleanSurvival, which integrate preprocessing for survival analysis, can make survival studies easier and quicker to perform, as well as make the results more robust.
 1 Introduction
@@
 However, it is also overlooked in the context of automated machine learning (AutoML) pipelines, which aim to streamline the ML development process by automating tasks such as algorithm selection, hyperparameter tuning and model evaluation—and increasingly incorporate data preparation as part of their pipelines.
-In this paper, we introduce CleanSurvival, an automated data preprocessing framework tailored for survival analysis.
+There exist several tools focusing on preprocessing for classification and regression models [27].
+In this paper, we introduce CleanSurvival, an automated data preprocessing framework tailored specifically for survival analysis.
 CleanSurvival leverages reinforcement learning (RL) techniques, specifically $Q$-learning, to optimize decisions such as imputation of missing values, detection and handling of outliers and feature extraction for survival models.
@@
 The article is organized as follows.
-Section 2 provides an overview of data preprocessing, survival analysis, AutoML and $Q$-learning.
-Section 3 reviews existing approaches to automated data preprocessing and AutoML frameworks.
+Section 2 provides an overview of data preprocessing, survival analysis, AutoML tools and $Q$-learning.
 Section 3 describes the architecture of CleanSurvival and its features.
-Section 5 presents the results of experimental evaluation of the framework on real-world datasets.
+Section 4 presents the results of experimental evaluation of the framework on real-world datasets and Section 5 discusses CleanSurvival in comparison to other upcoming tools in this field.
 Finally, Section 6 discusses the results and outlines future directions for research.
@@
 Data preprocessing involves cleaning, transforming and organizing raw data into a suitable format for analysis, and is a important step in the ML pipeline.
-Sub-tasks of data preprocessing include imputation or removal of missing values, detection and handling of outliers, variable selection and feature extraction and data transformations.
+Sub-tasks of data preprocessing include imputation or removal of missing values, detection and handling of outliers, variable selection, feature extraction and data transformations.
 These steps can have a profound downstream impact on classification performance [22] and model explanations [40].
@@
 In the latter case, one models survival—or not—in a predefined time window, reducing the problem to binary classification and losing granular information about event times [21].
-Survival analysis models, such as the Cox proportional hazards model, Kaplan–Meier estimator and accelerated failure time models, are widely used in practice.
+Survival analysis models, such as the Cox proportional hazards model, Kaplan–Meier estimator and accelerated failure time models, are widely used in practice [50, 41].
 These models estimate the hazard function, survival function, or survival probabilities over time, providing valuable insights into the relationship between covariates and survival outcomes.
-Concordance indices like the C-index are widely used in evaluation of survival models due to their simplicity and ease of interpretation [16].
+Concordance indices like the C-index are widely used in the evaluation of survival models due to their simplicity and ease of interpretation [23].
 The C-index evaluates the model’s ability to correctly rank individuals based on their risk of experiencing the event.
 A higher C-index indicates better discriminatory power.
-However, concordance indices have significant limitations, due to their poor calibration and failure to consider the distribution of survival times as well as their ranks.
-To measure how well survival probabilities align with observed event times, calibration metrics like Houwelingen’s $\alpha$ or D-calibration can be used [35] , and should be compared against a baseline model, such as the Kaplan–Meier estimator.
+However, concordance indices have significant limitations, due to their poor calibration and failure to consider the distribution of survival times as well as their ranks [16].
+The Brier Score is another metric in survival analysis which calculates the accuracies of the predicted survival probabilities at a given time point [42].
+Other calibration metrics like Houwelingen’s $\alpha$ or D-calibration can also be used to measure how well survival probabilities align with observed event times [49].
+They could be compared against a baseline model such as the Kaplan–Meier estimator.
 Weighted integrated survival log loss or integrated Graf score are recommended scoring rules [43].
@@
 $\text{IGS}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{k}\Delta t_{j}\,\frac{\left(S(t_{j}|x_{i})-\text{Observed}(t_{j},i)\right)^{2}}{G(t_{j})},$
-for a set of time points $\{t_{1},t_{2},\ldots,t_{k}\}$, where $\Delta t_{j}$ represents the difference between time points, $n$ is the number of individuals and $\text{Observed}(t,i)$ is an indicator function, equal to $1$ if the individual is known to have survived beyond $t$ and 00 otherwise.
+for a set of time points $\{t_{1},t_{2},\ldots,t_{k}\}$, where $\Delta t_{j}$ represents the difference between time points, $n$ is the number of individuals and $\text{Observed}(t,i)$ is an indicator function, equal to $1$ if the individual is known to have survived beyond $t$ and 0 otherwise.
 2.2 Automated machine learning
@@
 Mahdavi et al.
-[18] highlighted the potential of AI to solve data quality problems through data profiling and learning from past cleaning attempts [see also 17].
-A holistic approach integrates the cleaning process with downstream tasks so that the cleaning is optimized for predictive performance [20]; indeed when integrated into an AutoML framework some cleaning steps may be more important than others [21].
-Survival analysis in particular faces particular challenges, partly due to the relative lack of support for such models in the first place (versus classification or regression), as well as unique difficulties of handling censored time-to-event data [36], which are typically not addressed in conventional AutoML frameworks and do not feature in AutoML surveys [e.g.
-1].
+[25] highlighted the potential of AI to solve data quality problems through data profiling and learning from past cleaning attempts ([24].
+A holistic approach integrates the cleaning process with downstream tasks so that the cleaning is optimized for predictive performance [29]; indeed when integrated into an AutoML framework, some cleaning steps may be more important than others [30].
+Survival analysis in particular faces particular challenges, partly due to the relative lack of support for such models in the first place (versus classification or regression), as well as unique difficulties of handling censored time-to-event data [50], which are typically not addressed in conventional AutoML frameworks and do not feature in AutoML surveys [2].
 Seamlessly integrating these domain-specific preprocessing steps with the downstream tasks of model optimization and evaluation is a complex, underexplored area.
-Prominent AutoML pipelines include Auto-WEKA, an early AutoML system that uses Bayesian optimization to search for the best combination of preprocessing steps and machine learning algorithms [34]; TPOT, a tree-based pipeline optimization tool that uses genetic programming to evolve pipelines of data cleaning and machine learning operations [22]; and auto-sklearn, an extension of Auto-WEKA that incorporates more recent advancements in machine learning and hyperparameter optimization while offering a familiar interface based on the Python package scikit-learn [8].
-These AutoML pipelines have demonstrated promising results in various domains, including image classification, natural language processing and tabular data analysis.
-However, they often focus on general machine learning tasks and may not be specifically tailored to the challenges of survival analysis, particularly in the presence of missing data.
-Salhi et al.
+Prominent AutoML pipelines include Auto-WEKA, an early AutoML system that uses Bayesian optimization to search for the best combination of preprocessing steps and machine learning algorithms [48].
+It covers a range of models but does not cover crucial survival models like Cox or survival metrics [18].
+TPOT, a tree-based pipeline optimization tool that uses genetic programming to evolve pipelines of data cleaning and machine learning operations, however it performs simple imputation and does not perform censoring natively which means it cannot work on survival analysis [31].
+Similarly, auto-sklearn, an extension of Auto-WEKA that incorporates more recent advancements in machine learning and hyperparameter optimization while offering a familiar interface based on the Python package scikit-learn cannot perform preprocessing for survival models and metrics [13].
 [37] presented a recent survey of data preprocessing using AutoML (though survival analysis is not mentioned).
@@
 scikit-learn).
-Mumuni and Mumuni [19] also surveyed automated data processing for deep learning applications, highlighting in more detail the extent to which data processing steps are integral components of the automated pipeline.
+[27] also surveyed automated data processing for deep learning applications, highlighting in more detail the extent to which data processing steps are integral components of the automated pipeline.
 They similarly note the lack of early support from autosklearn.
+Table 1: Comparison of AutoML frameworks with respect to data preprocessing and survival analysis capabilities
+Framework License Missing values Categorical Imbalance Survival
+auto-sklearn Open-source Requires NaN removal Yes Yes External libraries
+TPOT Open-source Yes No No No
+Google AutoML Tables Proprietary Yes Yes Yes No
+H2O AutoML Open-source Yes Yes Yes Limited
+PyCaret Open-source Yes Yes Yes Limited
+AutoKeras Open-source Yes Yes Limited Custom models
+Azure AutoML Proprietary Yes Yes Yes External tools
+BigML Commercial Yes Yes Limited No
+MLflow Open-source No No No No
+FLAML Open-source No No Yes No
+MLJAR Open-source Yes Yes Limited No
+DataRobot Proprietary Yes Yes Yes Limited
+Amazon SageMaker Autopilot Proprietary Yes Yes Yes No
+AutoGluon Open-source Yes Yes Limited External libraries
+Table 1 compares the features of various AutoML solutions.
+Of frameworks offering automated data cleaning, DataRobot is proprietary, commercial platform and only appears to offer time-to-event modelling via a ‘hack’ of converting the task to a classification problem via discretization (see e.g.
+[9]).
+Google AutoML Tables, Amazon SageMaker Autopilot and Azure ML can perform different preprocessing steps but they are not always tailored to survival analysis and these are all proprietary tools [28].
+Big ML, that performs decision tree-based optimization, performs preprocessing but no survival analysis [12].
+MLJAR and AutoGluon both use stack ensembling, perform most of the preprocessing steps needed but do not suppost survival analysis [26, 46].
+Meanwhile, PyCaret can perform preprocessing and survival analysis as well but only to an extent.
+It can do mortality predictions but needs external tools for further survival analysis [35].
+Likewise, H2O.ai can run Cox proportional hazards models [8] as a fixed model, but not via its AutoML interface as it requies advanced platforms for the same.
+However, some dedicated data cleaning solutions have been proposed.
+[4] proposed Auto-Prep, an interactive Python-based tool that recommends data cleaning methods to the user based on application of candidate techniques and subsequent evaluation using simple classifiers or regression models.
+In a review of data preprocessing in AutoML (Section C) the authors highlight the capabilities, or lack thereof, of AutoML tools to perform data preprocessing and feature engineering without manual human intervention.
+Another Python package, Atlantic [38], automates preprocessing steps including feature engineering and missing value imputation for supervised learning tasks.
+The framework identifies the best combination of steps based on evaluation using tree-based model ensembles.
+MLsurvival [51] described an automated tool for cancer survival prediction that removes or imputes missing values, selects and standardizes features, trains survival models and then makes predictions.
+Unfortunately, neither a full text article nor open source implementation of the method were published.
+More recently, Pomsuwan and Freitas [34] proposed an AutoML system for survival analysis based on genetic algorithms and a combination of elastic-net Cox models, random survival forests and survival trees, optimizing for C-index.
+However, the tool does not incorporate data preprocessing.
 2.3 $Q$-learning
@@
 Other reinforcement learning approaches, including deep reinforcement learning [17] and Monte Carlo tree search [11], provide flexibility but introduce additional computational overhead and may require constraints or tailored mechanisms to ensure valid ML pipelines.
-3 Related work
-Table 1 compares the features of various AutoML solutions; as highlighted in Section 2, few offer integrated support for survival analysis and data preprocessing is not always within the optimization loop.
-Table 1: Comparison of AutoML frameworks and their native support for survival analysis and dynamic data preprocessing
-Framework Method Preprocessing Survival
-Amazon SageMaker Autopilot Bayesian optimization and ensembles Yes No
-auto-keras Neural architecture search Yes No
-auto-sklearn Bayesian optimization Limited No
-AutoGluon Stack ensembling Yes No
-Azure AutoML Bayesian optimization and meta-learning Yes No
-BigML Decision tree-based optimization Yes No
-DataRobot Proprietary ensemble and optimization Yes Limited
-FLAML Cost-aware Bayesian optimization No No
-Google AutoML Tables Neural architecture search Yes No
-H2O AutoML Random search and stacked ensembles Yes No
-MLflow Manual configuration Yes No
-MLJAR Random search and stacked ensembles Yes No
-PyCaret Iterative search with pipeline tuning Yes Limited
-TPOT Genetic programming No No
-Of frameworks offering automated data cleaning, DataRobot is proprietary, commercial platform and only appears to offer time-to-event modelling via a ‘hack’ of converting the task to a classification problem via discretization [see e.g.
-5].
-Meanwhile, H2O.ai can run Cox proportional hazards models as a fixed model, but not via its AutoML interface.
-However, some dedicated data cleaning solutions have been proposed.
-Bilal et al.
-[3] proposed Auto-Prep, an interactive Python-based tool that recommends data cleaning methods to the user based on application of candidate techniques and subsequent evaluation using simple classifiers or regression models.
-In a review of data preprocessing in AutoML (Section C) the authors highlight the capabilities, or lack thereof, of AutoML tools to perform data preprocessing and feature engineering without manual human intervention.
-Another Python package, Atlantic [28], automates preprocessing steps including feature engineering and missing value imputation for supervised learning tasks.
-The framework identifies the best combination of steps based on evaluation using tree-based model ensembles.
 Berti-Equille [3] developed Learn2Clean, a tool offering an innovative approach to data preprocessing.
@@
 It also has a complex dependency structure, which can make initial setup challenging for end-users.
-MLsurvival [37] described an automated tool for cancer survival prediction that removes or imputes missing values, selects and standardizes features, trains survival models and then makes predictions.
-Unfortunately, neither a full text article nor open source implementation of the method were published.
-More recently, Pomsuwan and Freitas [25] proposed an AutoML system for survival analysis based on genetic algorithms and a combination of elastic-net Cox models, random survival forests and survival trees, optimizing for C-index.
-However, the tool does not incorporate data preprocessing.
+We build upon Learn2Clean, by adapting the reinforcement learning framework to survival analysis.
 3 Methodology
@@
   • Simple mean/median imputation replaces missing values with the mean or median of the observed values for a given variable.
-  • Multiple imputation using chained equation [MICE, 4] provides a more robust approach by generating multiple suitable replacements for each missing value, creating several complete datasets for analysis.
+  • Multiple imputation using chained equation [5] provides a more robust approach by generating multiple suitable replacements for each missing value, creating several complete datasets for analysis.
 This method utilizes an iterative imputer, which starts with an initial guess and refines the estimates until convergence.
@@
 3.1.2 Outliers
-To ensure the reliability and validity of survival analysis, robust outlier detection methods are essential.
+To ensure the reliability and validity of survival analysis, robust outlier detection methods are essential [6].
   • For multivariate datasets, CleanSurvival employs the Elliptic Envelope algorithm [36] to identify and remove outliers based on their Mahalanobis distances from the data centre.
@@
 The tool generates random preprocessing pipelines and evaluates their performance, providing insights into the impact of different preprocessing choices.
-This mdoe can serve as a baseline for comparison with the optimized pipeline.
+This mode can serve as a baseline for comparison with the optimized pipeline.
 Custom pipeline
@@
 All methods were implemented in Python.
-Source code and documentation are available at https://github.com/datasciapps/CleanSurvival.
+Source code and documentation are available at https://github.com/phoenix0401/CleanSurvival.
 Details of the datasets used are provided in Table 2.
@@
 Rotterdam 2982 14 Rotterdam Study
-Flchain 7874 11 UCI Machine Learning Repository
+Flchain 7874 11 UCI Repository
 The results of data preprocessing strategies suggested by CleanSurvival are compared against the following methods:
   1. 1.
-    complete case analysis (CCA), effectively ignoring the problem of missingness and outliers
+    Complete case analysis (CCA): effectively ignoring the problem of missingness and outliers
   2. 2.
-    random selection of imputation methods, simulating a grid search over possible analysis choices
+    Random selection of imputation methods: simulating a grid search over possible analysis choices
   3. 3.
@@
 We evaluate each method based on quality metrics as well as the time taken to retrieve an acceptable solution.
-Hyperparameters that are not part of the CleanSurvival search space were left at default values, with architectural choices for the deep learning framework as follows: the DeepHit architecture consists of a shared sub-network with 2 hidden layers, each containing 100 neurons, using ReLU as the activation function.
+Hyperparameters that are not part of the CleanSurvival search space were left at default values.
+The architectural choices for the deep learning framework were as follows: the DeepHit architecture consists of a shared sub-network with 2 hidden layers, each containing 100 neurons, using ReLU as the activation function.
 This is followed by cause-specific sub-networks, each containing 2 hidden layers with 50 neurons per layer.
@@
 Since the Rotterdam dataset does not contain missing values, three variations of the dataset were used each representing a different missingness strategy (MCAR, MAR, MNAR) which were generated using the Jenga framework [39].
+After generating pipelines using CleanSurvival and Random, we compute Brier scores across timepoints (n=50) for the Rotterdam dataset, with three mechanisms and varying levels of missingness.
 4.2 Results
@@
 Flchain Existing 0.695 0.621 0.655
-(a) MCAR
-(b) MAR
-(c) MNAR
-(d) Flchain
+The C-index values in Table 3 show that CleanSurvival leads to better performance of the models across datasets with varying degrees of missingness.
+It has a higher C-index compared to CCA and Mean imputation.
+With Rotterdam datasets having high C-index and Flchain, this could be due to the different mechanisms of simulated missingness versus the natively-present missing values in FlChain.
+Refer to caption Refer to caption Refer to caption
+(a) MCAR (b) MAR (c) MNAR
+Figure 2: CleanSurvival vs Random
+Figure 2 shows the Brier scores over time for various missingness percentages as well as for the three mechanisms used to simulate missingness in the Rotterdam dataset.
+With timepoints chosen to be 50, it can be seen that overall CleanSurvival generated more stable Brier scores compared to the baseline of randomly generated pipelines.
+For the MCAR mechanism of missingness, across all levels of missingness, the Brier scores are stable and end at a lower score compared to Random.
+This is visible for both types of models - COX and RSF.
+For MAR and MNAR, apart from one missingness level (40 for the former and 30 for the latter), other line plots are all stable for the COX and RSF models when pipelines were preprocessed with CleanSurvival.
+These plots show that CleanSurvival can help improve robustness and consistency across different models and runs.
+Refer to caption Refer to caption Refer to caption Refer to caption
+(a) MCAR (b) MAR (c) MNAR (d) Flchain
 Figure 3: CleanSurvival vs Grid Search
@@
 These findings highlight the potential of CleanSurvival as a valuable tool for accelerating and automating the data preprocessing stage of survival analysis.
-By efficiently identifying optimal pipelines, not only does it save time and resources but it also enables researchers to focus on the core aspects of their analysis, ultimately leading to more robust and reliable results.
+By efficiently identifying optimal pipelines, not only does it save time and resources, but it also enables researchers to focus on the core aspects of their analysis, ultimately leading to more robust and reliable results.
+5 Discussion
+With most of the AutoML preprocessing tools focusing on classification and regression models, we present CleanSurvival as a tool which builds upon Learn2Clean to tailor preprocessing for survival analysis using a $Q$-learning-based framework and reward structures.
+New tools like AutoPrognosis [1] have developed automated tools using Bayesian Optimization that include survival analysis, but there is less scope for customisation of the pipelines.
+The widely used library scikit-survival integrates scikit-learn preprocessing with survival models but it requires manual pipeline building [7].
+The R library ‘mlr3proba’ provides a machine learning interface for survival analysis and preprocessing, however missingness is still dealt with externally [44].
+CleanSurvival can be performed on various survival models, including neural networks.
+It can handle different preprocessing steps required for survival analysis data.
+The user can also customise pipelines based on their requirement.
+This makes CleanSurvival an important tool in the emerging field of automation of machine learning and preprocessing pipelines.
 6 Conclusion
-In this paper, we have introduced CleanSurvival, the first ever $Q$-learning-based framework that automates data preprocessing for survival analysis, addressing the often underserved challenge of automation in presence of censored data.
+In this paper, we have introduced CleanSurvival, a $Q$-learning-based framework that automates data preprocessing for survival analysis, addressing the often underserved challenge of automation in the presence of censored data.
 We demonstrated its ability to enhance predictive performance and computational efficiency across different settings, compared to conventional methods or heuristic approaches.
@@
 Additionally, ensembling over differently cleaned datasets, together with raw data, may have the potential to increase predictive performance.
-7 Acknowledgement
+7 Declarations
+7.1 Ethics approval and consent to participate
+Not applicable. This study used publicly available datasets.
+7.2 Consent for publication
+Not applicable.
+7.3 Availability of data and materials
+All datasets and code are available at https://github.com/phoenix0401/CleanSurvival.
+7.4 Competing interests
+The authors declare no competing interests.
+7.5 Funding
+This work was supported by the German Federal Ministry of Education and Research via the project “Eventful” under grant ID 01IW23005.
+7.6 Authors’ contributions
+YK conceptualized the package, wrote and implemented the code and wrote the manuscript.
+DS provided supervision, assisted in conceptualizing, writing and editing manuscript.
+GG helped in code optimization and implementation.
+SV provided supervision, funding acquisition and helped in reviewing and editing the manuscript.
+KP performed data evaluation, code implementation and editing the manuscript.
+7.7 Acknowledgement
 We are grateful for the input of Sergey Redyuk, whose suggestions greatly improved the manuscript.
-This work was supported by the German Federal Ministry of Education and Research via the project “Eventful” under grant ID 01IW23005.